OpenStudy (anonymous):

Find the second derivative of the function. g(x) = 5sqrt[x] + e^(2x) ln(x) y"(x)=

OpenStudy (anonymous):

thanks for ur previous help:)i appreciate it

myininaya (myininaya):

\[g=5x^\frac{1}{2}+e^{2x} \ln(x)\] \[g'=5 \cdot \frac{1}{2} x^{\frac{1}{2}-1}+2e^{2x} \ln(x)+e^{2x} \cdot \frac{1}{x}\]

myininaya (myininaya):

\[g'=\frac{5}{2}x^\frac{-1}{2}+e^{2x}(2\ln(x)+\frac{1}{x})\] now we to do the derivative again \[g''=\frac{5}{2} \cdot \frac{-1}{2} x^{\frac{-1}{2}-1}+2e^{2x}(2 \ln(x)+\frac{1}{x})+e^{2x}(2 \frac{1}{x}-\frac{1}{x^2})\]

myininaya (myininaya):

np obsolete

OpenStudy (anonymous):

myininaya for step five with the second derivative shouldn't it be -5/4

myininaya (myininaya):

because 2 times 2 is 4 lol

myininaya (myininaya):

yes

myininaya (myininaya):

\[g''=\frac{-5}{4\sqrt{x^3}}+e^{2x}(4 \ln(x)+2\frac{1}{x}+2\frac{1}{x}-\frac{1}{x^2})\] \[g''=\frac{-5}{4\sqrt{x^3}}+e^{2x}(4 \ln(x)+\frac{4}{x}-\frac{1}{x^2}) \]

OpenStudy (zarkon):

myininaya your hands are going to fall off ;)

myininaya (myininaya):

lol

OpenStudy (anonymous):

yes sorry just pointing that out and also \[\sqrt{x ^{3}}= x \sqrt{x}\] haha @ Zarkon

OpenStudy (anonymous):

But that's just simplification

myininaya (myininaya):

i think i won't simplify anymore just because i forget to do stuff like multiply 2 times 2

OpenStudy (anonymous):

Yeah...just glad we caught that...algebra really sucks when you taking the derivative one small thing can mess u up lol

OpenStudy (anonymous):

hmm im really glad u guys r helping me out so what woud the the final answer be:S

OpenStudy (anonymous):

like if you ould just point out where u put it

OpenStudy (anonymous):

IT should be myininaya's fourth post after mine..In that post the second answer is the simplified one.

OpenStudy (anonymous):

I mean 3rd post

OpenStudy (anonymous):

hmm thanks so much, thats awfully long.r u sure its right?maybe this is one of the long oness

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Believe me I've down worse questions ones than this problem...it's right. What I saw was just small error. Of course when you are typing sometimes you make mistakes too.

OpenStudy (anonymous):

derivatives i just posted my question. im stuck in the middle of it .

OpenStudy (anonymous):

chaise can u please read my question i posted it already